Write an equation that expresses the following relationship: x varies directly as z and inversely as the difference between y and p. Then solve the equation for y.
Write an equation that expresses the given relationship. Use k as the constant of variation.
x=

Answer by stanbon(75887)

Write an equation that expresses the following relationship:
x varies directly as z and inversely as the difference between y and p.
x = k*z/(y-p)
-----------------------------
Then solve the equation for y.
y-p = kz/x
y = [kz/x]+p
===============
Cheers,
Stan H.
===============